OpenStudy (anonymous):

please help one question medal + fan *note: i am not asking for an answer but i am asking for help and guidance. i will admit i am not sure how to do this but I'm still not asking for an answer just patience and guidance.*

OpenStudy (anonymous):

(if i am gone its because i am helping someone but no worry i still will come here and work)

OpenStudy (cwrw238):

You can find the centre and radius by converting to the standard form (x - a)^2 + ( y - b)^2 = r^2 where the centre is (a,b) and r = the radius.

OpenStudy (cwrw238):

you can do this by completing the square on the terms in x and the terms in y: x^2 + 4x + y^2 - 6y= -4 (x + 2)^2 -4 + y^2 - 6y = -4 - thats completing the square on the terms in x. You do the same thing with y^2 + 6x

OpenStudy (cwrw238):

to complete the square use the identity x^2 + bx = (x +(b/2)^2 - (b/2)^2

OpenStudy (cwrw238):

move all the numbers to RHS of the '=' . this number will equal r^2

OpenStudy (cwrw238):

* error back 3 posts that y^2 - 6y not y^2 + 6x

OpenStudy (anonymous):

O.O I'm sorry, I'm a little confused. @cwrw238

OpenStudy (anonymous):

@jdoe0001 @jim_thompson5910 @superdavesuper

jimthompson5910 (jim_thompson5910):

What's the general equation for a circle? Do you have that in your notes?

OpenStudy (anonymous):

Yes I do, please give me a sec to type it in.

jimthompson5910 (jim_thompson5910):

ok take your time

OpenStudy (anonymous):

(x–h)2 + (y–k)2 = r2, where (h, k) is the center and r is the radius.

jimthompson5910 (jim_thompson5910):

I think you meant to write (x-h)^2 + (y-k)^2 = r^2 right?

OpenStudy (anonymous):

oh oops lol yeah forgot to put the "^"

jimthompson5910 (jim_thompson5910):

ok so we have this to start with x^2 + 4x + y^2 - 6y= -4

jimthompson5910 (jim_thompson5910):

is that in the form (x-h)^2 + (y-k)^2 = r^2 ?

OpenStudy (anonymous):

doesn't look like it.

jimthompson5910 (jim_thompson5910):

correct, it is not in that form

jimthompson5910 (jim_thompson5910):

so we have to transform it into that form

jimthompson5910 (jim_thompson5910):

to do that we complete the square for the x and y terms

jimthompson5910 (jim_thompson5910):

focus on just the x terms (either with an x or x^2 in them) x^2 + 4x is this a perfect square? if so, then what does it factor to? if not, then what is missing to make it a perfect square?

OpenStudy (anonymous):

I think it is a perfect square?

jimthompson5910 (jim_thompson5910):

what does x^2 + 4x factor to? a perfect square is something in the form S^2 where S is some expression example: x^2 is a perfect square, so is (y-10)^2, so is 225 (since 225 = 15^2)

jimthompson5910 (jim_thompson5910):

so are you able to put x^2 + 4x into the form S^2 ?

OpenStudy (anonymous):

@jim_thompson5910 im so sorry for asking for your help. i literally just went through my lesson again and there is a step-by-step on solving this, i am truly sorry. i will still give you a medal and i am already fanned to you. thanks so much.

jimthompson5910 (jim_thompson5910):

that's ok. I'm glad you found the solution to it